OpenStudy (anonymous):

The graph of which of the following equations contains the points (2, 3) and (3, 2)? y = x - 1 y = 5 - x y = 7 - 2x

OpenStudy (anonymous):

See, you are given points, just substitute in the equations given, and check which one is satisfying those points..

OpenStudy (anonymous):

In \((2,3)\), \(x=2\) and \(y=3\).. So, put these values in, for example, first equation: \(y = x-1 \\ 3 = 2-1 \\ but \; 3 \ne 1\), so this point is not satisfying first equation.. Similarly, check for other two equations..